OpenStudy (anonymous):

Find the height of a shorter building given the following: the height of the taller building, the angle of depression between the two buildings, and the distance between the buildings. How would I set this problem up?

OpenStudy (anonymous):

 So \[\tan(\theta) = y /x --> x * \tan(\theta) = y\] h2 = h1 - y I may have the wrong angle. wasn't sure what angle of depression was.